OpenStudy (anonymous):

I need help with inverses, what is the inverse of f if f(x)=3sqrt(x-2)

OpenStudy (campbell_st):

just swap x and y\[x = 3\sqrt{y - 2}\] and now make y the subject

OpenStudy (anonymous):

f-1(x)=2+ In(x)/In(3)

OpenStudy (campbell_st):

ingore the posted solution from @loveyourlife.... it's wrong

OpenStudy (michele_laino):

please, keep in mind that for existence of you function has to be: \[x \ge 2\], after that we can write: if y=3sqrt(x-2), then

OpenStudy (anonymous):

Okay than I am sorry if it is wrong but I tried to help that's all sorry again !

OpenStudy (michele_laino):

\[x=f ^{-1}(y)=2+\frac{ y ^{2} }{ 9 }\]

OpenStudy (anonymous):

So that would make the answer ?

OpenStudy (michele_laino):

@loveyourlife you are welcome!!

OpenStudy (campbell_st):

well go and try to follow the methods provided... Open Study is about helping understanding and not giving answers

OpenStudy (anonymous):

@michele_laino if he's wrong what's right?

OpenStudy (michele_laino):

since @campbell_st is right!, then I will go to explain my answer: now f(x) is: \[y=3\sqrt{x-2}\] please @Claydirt square both sides of that equation

OpenStudy (michele_laino):

please try, you have to perform this calculus: \[(y)^{2}=(3\sqrt{x-2})^{2}\] please continue @Claydirt

OpenStudy (anonymous):

I got f^-1(x)=x^3+2 is that correct ?

OpenStudy (michele_laino):

No, I think it is a wrong answer, please try to perform the calculus above that I began

OpenStudy (michele_laino):

left side is equals to: \[y ^{2}\] and right side is equals to: \[9*(...)\] please continue

OpenStudy (michele_laino):

it is equals to \[9(x-2)\] so: \[y ^{2}=9(x-2)\] please solve that equation in order to find x, and you will get your answer